The body mass index,I, can be used to determine an individuals risk for heart disease. An index less than 25 indicates a low risk. The body mass index is given by the formula, or model, I=(700w/H^2), where W=weight, in pounds, and H=height, in inches. Francis wieghs 216 pounds and stands 74 inches tall. What is his approcimate body mass index? Find an ineqyality describing all weights W that Francis can have and be in the low-risk category.
Answer by mananth(16949)

The body mass index,I, can be used to determine an individuals risk for heart disease. An index less than 25 indicates a low risk. The body mass index is given by the formula, or model, I=(700w/H^2), where W=weight, in pounds, and H=height, in inches. Francis wieghs 216 pounds and stands 74 inches tall. What is his approcimate body mass index? Find an ineqyality describing all weights W that Francis can have and be in the low-risk category.
w=216
h=74
I= (700*216)/(74)^2
I = 27.61
...
25= (700*w)/(74)^2
25*(74)^2=700w
(25*974)^2/700 = w
196 pounds
{xeR|x<196}
...
m.ananth@hotmail.ca